# MUSCLE: A Model Update Strategy for Compatible LLM Evolution

### v1

@@ Abstract
Large Language Models (LLMs) are frequently updated due to data or architecture changes to improve their performance. When updating models, developers often focus on increasing overall performance metrics with less emphasis on being compatible with previous model versions. However, users often build a mental model Bansal et al. (2019) of the functionality and capabilities of a particular ML model they are interacting with. They have to adapt their mental model with every update – a draining task that can lead to user dissatisfaction. In practice, fine-tuned downstream task adapters rely on pretrained LLM base models. When these base models are updated, these user-facing downstream task models experience instance regression or negative flips –previously correct instances are now predicted incorrectly. This happens even when the downstream task training procedures remain identical. Our work aims to provide seamless model updates to a user in two ways. First, we provide evaluation metrics for a notion of compatibility to prior model versions, specifically for generative tasks but also applicable for discriminative tasks. We observe regression and inconsistencies between different model versions on a diverse set of tasks and model updates. Second, we propose a training strategy to minimize the number of inconsistencies in model updates, involving training of a compatibility model that can enhance task fine-tuned language models. We reduce negative flips – instances where a prior model version was correct, but a new model incorrect – by up to 40% from Llama 1 to Llama 2.
@@
Jessica Echterhoffthanks:   Research conducted during internship at Apple. Corresponding authors: jechterh@ucsd.edu,    mpouransari@apple.com Fartash Faghri Raviteja Vemulapalli UC San Diego Apple Apple Ting-Yao Hu Chun-Liang Li Oncel Tuzel Hadi Pouransari Apple Apple Apple Apple
@@ 1 Introduction
Large Language Models (LLMs) are often pretrained on large-scale corpora to obtain a base model that has general world knowledge. These base models are typically evaluated using a suite of benchmarks that mostly focus on zero/few-shot performance and in-context learning capabilities. Training these models is expensive and only a few organizations have access to the resources needed for this. Hence, to enable various user-facing applications such as summarization, chatbots, code assistants, and question answering, practitioners often adapt pretrained base models by training task-specific parameter-efficient adapters using downstream task datasets.
@@
Figure 1: A real example of a model update that introduces negative flips, where previously correct predictions become incorrect. With our model update strategy using a compatibility adapter approach, we can enhance compatibility with the previous model while maintaining the performance gain (measured by the ROUGE-1 score for the summarization task in this figure) of the model update.
@@
Several scenarios drive updates to the base model. For example, improved training strategy, advances in LLM architectures or increasing model context length Touvron et al. (2023a), the availability of additional or higher quality datasets Gunasekar et al. (2023), the expansion of model vocabulary (e.g., to support multilingual or multimodal models), or simply training for a longer period  Biderman et al. (2023). When a base model is updated, the associated task adapters need to be retrained to have meaningful downstream models. Hence, in the rest of the paper, we use the term model update to refer to an update to a downstream task model, which includes updating the base model and retraining the task adapter.
When a model is updated, each data sample can be categorized into one of the four groups depicted by four quadrants shown in Fig. 2. The new model could result in a negative flip (quadrant 4) for many samples (Yan et al., 2021) even when it has a better overall performance when compared to the previous model. A real example of a negative flip caused by a model update is shown in Fig. 1 for diagloug summarizatoin task. This regression behavior could confuse the user driving down the user satisfaction (Sakai, 2022). Furthermore, the importance of regression testing has been recently studied for evolving LLMs used through APIs (Ma et al., 2023).
@@
Many existing works focus on increasing positive flips (quadrant 2) while reducing negative flips (quadrant 4) Cai et al. (2022); Sakai (2022); Yan et al. (2021); Li et al. (2023b); Schumann et al. (2023). However, yhey neglect prediction inconsistencies in the scenarios where both model versions are incorrect (quadrant 3). For example, Träuble et al. (2021) assumes the cost of flips from one incorrect class to another incorrect class to be zero. We argue that there is value in being consistent when both models are incorrect. A user may have developed coping strategies on how to interact with a model when it is incorrect; therefore, inconsistencies in mistakes can lead to user dissatisfaction.
@@
Previous works have mainly addressed the regression challenge for classification tasks Cai et al. (2022); Sakai (2022); Yan et al. (2021); Li et al. (2023b); Schumann et al. (2023). In this work, we systematically study the problem of downstream model compatibility (reducing negative flips and being consistent when making mistakes) using various downstream tasks (both discriminative and generative) and different base models.
@@
  • We formulate the compatibility problem when updating LLMs. We consider a setup where base LLMs are updated independently and study model compatibility and performance through parameter-efficient task-specific fine-tuning.
  • We extend notion of model compatibility from discriminative to generative tasks, and propose regression/compatibility metrics that consider similarities in model behavior after an update, going beyond the negative flip rate metric.
  • We investigate model compatibility for different update scenarios using open-weights models and find significant regression across various tasks.
  • We propose learning a compatibility adapter to align model versions and minimize regression. We demonstrate up to 40% reduction in negative flip rate (e.g. for Llama 1 to Llama 2 update in language understanding) and reduction in model inconsistency for other downstream tasks such as summarization, math reasoning, and commonsense question-answering.
Figure 2: Four possibilities arise for each sample when a model is updated. Quadrants 2 and 4 show positive and negative flips, respectively. A model update with more instances in quadrant 2 and fewer in quadrant 4 indicates more performance gain and less regression. Furthermore, quadrant 3 corresponds to instances where both models are incorrect. Encouraging similarity between the old and new models in this case (i.e., making the same mistakes) results in a more seamless model update from the user’s perspective.
@@ 2 Related Work
2.1 Measuring Regression Errors
@@
Yan et al. (2021) introduce the negative flip rate (NFR) to evaluate model compatibility for classification tasks. NFR calculates the fraction of instances that were previously correct but are now incorrect with the new model. Other metrics include backward trust compatibility (BTC) (Srivastava et al., 2020), the ratio of instances that the new model predicts correctly among all instances the old model predicts correctly. Ultimately, these metrics have been shown to measure similar statistics Sakai (2022). Matsuno and Sakuma (2023) propose backward compatibility with a conditional distribution, which computes the ratio at which the accuracy of the conditional distribution of the new model is equal to or higher than that of the old model. Cai et al. (2022) introduce the negative flip impact for graph NLP tasks, taking into account the negative flips and the overall error rate of the model. All of these metrics are limited to the evaliation of discriminative classification tasks.
2.2 Reducing Update Regression
@@
Prior work found that model ensembles reduce model update regression. This can be attributed to the reduction of variance by ensembling, as every single model may capture the training data from a distinct aspect (Yan et al., 2021; Xie et al., 2021). Other approaches include choosing the most centric model from the ensemble Xie et al. (2021), aligning two models’ uncertainties Li et al. (2023b), or using gating mechanisms Lai et al. (2023). To ensure model compatibility, previous work has also used model parts from the old model to infuse into the new one (Ran et al., 2023), with the limitation of both models being required at inference time. For limited use cases, Qin et al. (2023) have shown to re-use previously learned adapters when purely a data update was performed. These methods either introduce larger memory footprint by re-using old model parts, or are limited to (same-domain) data-updates and same models only.
@@
Originally proposed for model compression Buciluǎ et al. (2006), a (smaller) student model is trained to mimic a (larger) teacher model. By treating the old model as the teacher and the new model as the student, knowledge distillation has been shown to reduce model update regression in vision and language discriminative tasks Yan et al. (2021); Xie et al. (2021); Schumann et al. (2023); Jaeckle et al. (2023); Zhang et al. (2021); Shen et al. (2020); Ramanujan et al. (2022). Shen et al. (2020) propose a distillation-based influence loss to align new model representations with those of the old model. Similarly, Ramanujan et al. (2022); Jaeckle et al. (2023) apply distillation after a learned transformation module. Schumann et al. (2023) propose weight interpolation between the old and the new model, Zhao et al. (2022) suggest matching old and new model distributions, Yan et al. (2021) distill from an ensemble, and Caciolai et al. (2023) recommend using focal distillation. However, none of these approaches evaluate their approaches on generative tasks or require both the old and new models to be in memory.
@@ 3 Problem Formulation
Training LLMs from scratch for every downstream task is expensive. Hence, we follow the common setup of finetuning a pre-trained base LLM to multiple downstream tasks with task-specific LoRA adapters (Hu et al., 2021).
Let $\mathcal{M}^{\text{base}}_{i}$ denote the $i^{th}$ version of the base LLM with parameters $\theta_{i}$. We adapt $\mathcal{M}^{\text{base}}_{i}$ to a downstream task $\mathcal{T}$ using an adapter $\mathcal{A}_{i}^{\mathcal{T}}$ to obtain a downstream model $\mathcal{M}_{i}^{\mathcal{T}}$ with weights $\theta^{\mathcal{T}}_{i}=\theta_{i}+\Delta_{i}^{\mathcal{T}}$, where $\Delta_{i}^{\mathcal{T}}$ denotes the weights of the task-specific adapter $\mathcal{A}_{i}^{\mathcal{T}}$ learned using the training data corresponding to task $\mathcal{T}$. When the base model is updated from $\mathcal{M}^{\text{base}}_{v1}$ to $\mathcal{M}^{\text{base}}_{v2}$, the task-specific adapters are re-trained for each downstream task. Hereafter, for simplicity of notation, we use $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ to refer to the task-adapted models $\mathcal{M}_{v1}^{\mathcal{T}}$ and $\mathcal{M}_{v2}^{\mathcal{T}}$, respectively, and explicitly mention the task $\mathcal{T}$ when needed.
Figure 3: When updating a model, regression on individual tokens and instances can arise. We use a masked approach to select tokens to be aligned with knowledge distillation either with the old version to remain consistent, or with the new task model to increase performance.
@@ 3.1 Backward Compatibility Metrics
Here $\mathbf{1}$ denotes the indicator function. This notion of regression is partly applicable for autoregressively trained tasks. LLM benchmarks Gao et al. (2023) calculate the likelihood of every possible choice in a multiple-choice question and choose the response with the highest likelihood. This evaluation can be used to indicate regression, similar to classification for tasks with multiple choices Zellers et al. (2019); Wang et al. (2019); Welbl et al. (2017).
@@
Inconsistencies do not only occur within positive or negative flips. For example, when the old model predicts class A, the new model class B, and the ground truth is class C. Similarly, if there are multiple ground truth options, a flip could occur within the ground truth options, as in quadrant 1 in Fig. 2. We argue that if we cannot produce a better prediction, we should at least stay consistent to the users expectations and propose metrics extended negative flips metrics for this use case.
@@
In generative tasks for language models, we do not necessarily have multiple choices for which we can predict if an instance regressed or not. Hence a metric incorporating multiple choices and then calculating negative flips is not applicable for continuous evaluation metrics. For example, for summarization tasks, which continuous metrics such as ROUGE (Lin, 2004) or BERT Score (Zhang et al., 2019) scores are often used for evaluation. Hence, we require a new metric to evaluate compatibility for generative tasks like summarization.
@@ 3.2 Extended Evaluation Metrics
To extend the possibility of evaluating model updates, we propose a suite of metrics that evaluate model updates specifically for generative tasks that may or may not have multiple options available (e.g summarization tasks), as well as extend measurement of inconsistent flips.
Accounting for Flips within Gain or Regression
@@
which includes the possibility that neither of the models give the correct answer, but a change in behavior occurs. Similarly, for multi-label tasks this notion can account for ground-truths that may have flipped during a model update when multiple grount truth options exist.
Smooth Gain and Regression
To add fine-grained and continuous metrics for generative tasks like summarization, we evaluate the expected performance gain and regression compared to the previous version, to enable a user-centric model update in accordance to the four quadrants of Fig. 4. Our general framework aims to be independent of the actual similarity metric, such that it can be chosen in accordance with any respective task of interest.
@@
acting as an indication of the distance of the two model outputs with respect to the grand truth. This gain per instance enables us to classify which model update compatibility quadrant each instance falls into. In practice, similarity metrics could be BERT Score Zhang et al. (2019), ROUGE Score Lin (2004), BLEU Score Papineni et al. (2002), or model-as-a-judge metrics Huang et al. (2024) depending on the use case and task.
To get an indication similar to positive and negative flip rate in generative tasks, we observe the distribution of instances with positive gain or negative regression:
@@
To observe the magnitude of regression and gain, we can observe the expectation:
@@
This gives an indication of the average magnitude of change in similarity metric when gain or regression occurs.
@@
Dataset Task
HellaSwag (Zellers et al., 2019)
@@
SAMsum (Gliwa et al., 2019) Dialogue Summarization
Table 2: We tackle compatible model updates for different downstream tasks and datasets, including multiple-choice and generative tasks.

### v2

@@ Abstract
Large Language Models (LLMs) are regularly updated to enhance performance, typically through changes in data or architecture. Within the update process, developers often prioritize improving overall performance metrics, paying less attention to maintaining compatibility with earlier model versions. Instance-level degradation (instance regression) of performance from one model version to the next can interfere with a user’s mental model Bansal et al. (2019) of the capabilities of a particular language model. Users having to adapt their mental model with every update can lead to dissatisfaction, especially when the new model has degraded compared to a prior version for a known use case (model update regression). We find that when pretrained LLM base models are updated, fine-tuned user-facing downstream task adapters experience negative flips – previously correct instances are now predicted incorrectly. We observe model update regression between different model versions on a diverse set of tasks and models, even when the downstream task training procedures remain identical. We argue for the importance of maintaining model update compatibility during updates, and present evaluation metrics designed specifically for generative tasks, while also being applicable to discriminative tasks. We propose a training strategy to minimize the extent of instance regression in model updates, involving training of a compatibility adapter that can enhance task fine-tuned language models. We show negative flips reduce by up to 40% e.g. when updating Llama 1 to Llama 2 with our proposed method.
@@
Jessica Echterhoffthanks:   Research conducted during an internship at Apple. Corresponding authors: jechterh@ucsd.edu,    mpouransari@apple.com Fartash Faghri Raviteja Vemulapalli UC San Diego Apple Apple Ting-Yao Hu Chun-Liang Li Oncel Tuzel Hadi Pouransari Apple Apple Apple Apple
@@ 1 Introduction
Large Language Models (LLMs) are often pre-trained on large-scale corpora to obtain a base model with general world knowledge. These base models are typically evaluated using a suite of benchmarks that mostly focus on zero/few-shot performance and in-context learning capabilities. Training these models is expensive, and only a few organizations have access to the resources needed. Hence, to enable various user-facing applications such as summarization, chatbots, code assistants, and question-answering, practitioners often adapt pre-trained base models by training task-specific parameter-efficient adapters using downstream task datasets.
@@
Figure 1: A real example of a model update that introduces instance regression (negative flip, where a previously correct prediction becomes incorrect) (top). With our model update strategy using a compatibility adapter approach, we enhance model update compatibility to the previous model while maintaining the overall performance gain (e.g. measured by the ROUGE-1 score for the summarization task) of the model update (bottom).
@@
Several scenarios drive updates to the base model, e.g. improved training strategy, advances in LLM architectures or increasing model context length Touvron et al. (2023a), the availability of additional or higher quality datasets Gunasekar et al. (2023), the expansion of model vocabulary (e.g., to support multilingual or multimodal models), or simply training for a longer period  Biderman et al. (2023).
When a base model is updated, all associated task adapters need to be retrained/updated to have meaningful downstream models. Hence, in the rest of the paper, we use the term model update to refer to an update to a downstream task model, which includes updating the base model and retraining the task adapter.
Instance Regression An instance $i$ is predicted correctly by an old model $\mathcal{M}_{v1}$, but incorrectly by a new model $\mathcal{M}_{v2}$.
Model Update Regression Aggregation of all instances correctly predicted by an old model $\mathcal{M}_{v1}$, but incorrectly by an updated model version $\mathcal{M}_{v2}$.
Model Update Compatibility Measures the degree to which predictions made by an updated model $\mathcal{M}_{v2}$ remain consistent with the predictions of the previous model version $\mathcal{M}_{v1}$, specifically on instances where $\mathcal{M}_{v1}$ was correct. Quantifies the alignment or compatibility between model versions across updates.
Table 1: Definitions of Instance Regression, Model Update Regression, and Model Update Compatibility.
@@
When a model is updated, we evaluate model update compatibility with different metrics by four quadrants shown in Fig. 2. The new model could produce a worse prediction (negative flip (quadrant 4)) for many samples (Yan et al., 2021) even when it has a better overall performance when compared to the previous model. A real example of instance regression measured as a negative flip is shown in Fig. 1 for a dialogue summarization task. This kind of regression can confuse the user and impair their satisfaction (Sakai, 2022). We denote the aggregated overall regression for all individual instances as model update regression (Table 1). Regression testing has become increasingly important for the evolving use of LLMs accessed via APIs (Ma et al., 2023). When updating models, practitioners typically focus on increasing positive flips (quadrant 2) while avoiding negative flips (quadrant 4) Cai et al. (2022); Sakai (2022); Yan et al. (2021); Li et al. (2023b); Schumann et al. (2023). However, they neglect prediction inconsistencies in the scenarios where both model versions are incorrect (quadrant 3) or already correct, but slightly different (quadrant 1). For example, Träuble et al. (2021) assumes the cost of flips from one incorrect class to another incorrect class to be zero. We argue that there is value in evaluating consistency when both models are wrong. A user may have developed coping strategies on how to interact with a model when it is incorrect; therefore, inconsistencies in mistakes can lead to user dissatisfaction.
@@
Previous works have mainly addressed the model update regression challenge for classification tasks Cai et al. (2022); Sakai (2022); Yan et al. (2021); Li et al. (2023b); Schumann et al. (2023). In this work, we systematically study the problem of model update compatibility using discriminative and generative downstream tasks and different base models.
@@
  • We formulate the compatibility problem when updating LLMs. We focus on common setups where base LLMs undergo updates and evaluate model compatibility and performance via parameter-efficient, task-specific fine-tuning, as these models are deployed for user interaction.
  • We extend the notion of model compatibility from discriminative to generative tasks, and propose compatibility metrics that consider similarities in model behavior after an update, going beyond the negative flip rate metric.
  • We investigate model compatibility for different update scenarios using open-weight models and find significant model update regression across various tasks.
  • We propose learning a compatibility adapter to align model versions and minimize model update regression. We demonstrate up to 40% reduction in negative flip rate (e.g. for Llama 1 to Llama 2 update in language understanding) and reduced model inconsistency for downstream tasks such as summarization, math reasoning, and commonsense question-answering.
Figure 2: Four possibilities arise for each sample when a model is updated. Quadrants 2 and 4 show positive and negative flips, respectively. Quadrant 3 corresponds to instances where both models are incorrect. Encouraging similarity between the old and new models in this case (i.e., making the same mistakes) results in a more seamless model update from the user’s perspective.
@@ 2 Related Work
2.1 Measuring Model Update Regression
@@
Yan et al. (2021) introduce negative flip rate (NFR) to evaluate model compatibility for classification tasks. NFR calculates the fraction of instances that were previously correct but are now incorrect with the new model. A similar statistic Sakai (2022), backward trust compatibility (BTC) (Srivastava et al., 2020), measures the ratio of instances that the new model predicts correctly among all instances the old model predicts correctly. Matsuno and Sakuma (2023) propose backward compatibility with a conditional distribution, which computes the ratio at which the accuracy of the conditional distribution of the new model is equal to or higher than that of the old model. Cai et al. (2022) introduce the negative flip impact for graph NLP tasks, taking into account the negative flips and the overall error rate of the model. These aforementioned metrics are limited to the evaluation of discriminative classification tasks.
2.2 Reducing Model Update Regression
@@
Prior work found that model ensembles reduce model update regression. This can be attributed to the reduction of variance by ensembling, as every single model may capture the training data from a distinct aspect (Yan et al., 2021; Xie et al., 2021). Extensions on this line of work include choosing the most centric model from an ensemble Xie et al. (2021), aligning two models’ uncertainties Li et al. (2023b), or using gating mechanisms Lai et al. (2023). Previous work has also used model parts from an old model to infuse into the new one (Ran et al., 2023), with the limitation of both models being required at inference time. For limited use cases, Qin et al. (2023) have shown to re-use previously learned adapters when purely a data update was performed. All these methods either introduce a larger memory footprint by re-using old model parts or are limited to (same-domain) data-updates and same models.
@@
Originally proposed for model compression Buciluǎ et al. (2006), a (smaller) student model is trained to mimic a (larger) teacher model. By treating the old model as the teacher and the new model as the student, knowledge distillation has been shown to reduce model update regression in vision and language discriminative tasks Yan et al. (2021); Xie et al. (2021); Schumann et al. (2023); Jaeckle et al. (2023); Zhang et al. (2021); Shen et al. (2020); Ramanujan et al. (2022). Shen et al. (2020) propose a distillation-based influence loss to align new model representations with those of the old model. Similarly, Ramanujan et al. (2022); Jaeckle et al. (2023) apply distillation after a learned transformation module. Schumann et al. (2023) propose weight interpolation between the old and the new model, Zhao et al. (2022) suggest matching old and new model distributions, Yan et al. (2021) distill from an ensemble, and Caciolai et al. (2023) recommend using focal distillation. However, none of these approaches evaluate their approaches on generative tasks or they require both the old and new models to be in memory at inference time.
@@ 3 Problem Formulation
While existing methods provide valuable approaches to mitigating model update regression, they predominantly focus on discriminative classification tasks and often require additional memory at inference time. We propose a flexible approach to updating models without sacrificing performance or compatibility across downstream tasks.
@@
We follow the common setup of finetuning a pre-trained base LLM to multiple downstream tasks with task-specific LoRA adapters (Hu et al., 2021). Let $\mathcal{M}^{\text{base}}_{i}$ denote the $i^{th}$ version of a base LLM with parameters $\theta_{i}$. We adapt $\mathcal{M}^{\text{base}}_{i}$ to a downstream task $\mathcal{T}$ using an adapter $\mathcal{A}_{i}^{\mathcal{T}}$ to obtain a downstream model $\mathcal{M}_{i}^{\mathcal{T}}$ with weights $\theta^{\mathcal{T}}_{i}=\theta_{i}+\Delta_{i}^{\mathcal{T}}$, where $\Delta_{i}^{\mathcal{T}}$ denotes the weights of the task-specific adapter $\mathcal{A}_{i}^{\mathcal{T}}$ learned using the training data corresponding to task $\mathcal{T}$. When the base model is updated from $\mathcal{M}^{\text{base}}_{v1}$ to $\mathcal{M}^{\text{base}}_{v2}$, the task-specific adapters are re-trained for each downstream task. Hereafter, for simplicity of notation, we use $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ to refer to the task-adapted models $\mathcal{M}_{v1}^{\mathcal{T}}$ and $\mathcal{M}_{v2}^{\mathcal{T}}$, respectively, and explicitly mention the task $\mathcal{T}$ when needed.
@@ 3.1 Backward Compatibility Metrics
Here $\mathbf{1}$ denotes the indicator function. This notion of regression is partly applicable to autoregressively trained tasks. LLM benchmarks Gao et al. (2023) calculate the likelihood of every possible choice in a multiple-choice question and choose the response with the highest likelihood to calculate a likelihood based accuracy. This evaluation can indicate model update regression similar to classification for tasks when multiple choices are available Zellers et al. (2019); Wang et al. (2019); Welbl et al. (2017).
@@
Other inconsistencies arise when the old model predicts class A, the new model class B, and the ground truth is class C. Similarly, if there are multiple ground truth options, a flip could occur within the ground truth options, as in quadrant 1 in Fig. 2. These inconsistencies are not captured by positive or negative flips. We argue that if we cannot produce a better prediction, we should at least stay consistent with the user’s expectations and propose extended negative flips metrics for this use case.
Figure 3: When updating a model, regression on individual tokens and instances can arise. We use a masked approach to select tokens to be aligned with knowledge distillation either with the old version to remain consistent or with the new task model to increase performance.
@@
In generative tasks for language models, we do not necessarily have multiple choices for which we can predict if an instance regressed or not. Hence a metric incorporating multiple choices and then calculating negative flips is not applicable for continuous evaluation metrics. Typical continuous metrics in generative language tasks are ROUGE (Lin, 2004) or BERT (Zhang et al., 2019) scores. As regular negative flips are incapable of capturing these nuances, we require a new metric to evaluate compatibility for generative tasks like summarization.
@@ 3.2 Extended Evaluation Metrics
We propose a suite of metrics that evaluate model update compatibility on a fine-grained basis specifically for generative tasks (e.g summarization).
Accounting for Flips when Both Models are Incorrect
@@
This includes the possibility that neither of the models gives the correct answer, but a change in behavior occurs that can confuse a user. Similarly, for multi-label tasks, this notion can account for ground-truths that may have flipped during a model update when multiple ground truth options exist.
Smooth Compatibility Metrics
To add continuous metrics for generative tasks, we evaluate the expected model update regression. Our general framework aims to be independent of the actual similarity metric, such that it can be chosen in accordance with any respective task of interest. For example, translation tasks might require BLEU (Papineni et al., 2002), but summarization ROUGE (Lin, 2004) evaluation. Additionally, we want to measure a notion of performance gain, when both models are correct, but one might still be better than the other.
@@
acting as an indication of the distance between the two model outputs with respect to the grand truth. This per instance indication of distance enables us to classify which model update compatibility quadrant an instance falls into. In practice, similarity metrics could be BERT Score Zhang et al. (2019), ROUGE Score Lin (2004), BLEU Score Papineni et al. (2002), or model-as-a-judge metrics Huang et al. (2024) depending on the use case and task.
To get a metric for generative tasks that is similar to the positive and negative flip rate in classification tasks, we observe the distribution of instances with positive gain or negative regression:
@@
To observe the magnitude of model update regression and gain, we can observe the expectation:
@@
These give an indication of the average magnitude of change in a similarity metric when gain or regression occurs.
@@
Dataset Task Metric
HellaSwag
@@
Exact Match
Accuracy
SAMsum Dialogue Summarization ROUGE-1
Table 3: We tackle compatible model updates for different downstream tasks and datasets, including multiple-choice and generative tasks (Zellers et al., 2019; Bisk et al., 2019; Cobbe et al., 2021; Gliwa et al., 2019).